OpenStudy (anonymous):

Can someone help? Find the surface area of a conical grain storage tank with a height of 45 meters and a diameter of 20 meters?

OpenStudy (anonymous):

the formula for surface area of a cone is pi*r*l where l= slope length the height given is I think the vertical height so we have a triangle base 10 altitude 45 so get the hypotenuse= sqrt(10^2+45^2)= 46.1 surface area = pi*10*46.1= 461*pi =1448.27